OpenStudy (anonymous):

Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 (2, 4) and (1, 1) Set #2 (4, 6) and (2, 7) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (mertsj):

Do you know the slope formula?

OpenStudy (anonymous):

The slope is the rate of change. It can be calculated as the change in y values divided by the change in x values. Set 1 Change in y values = 4 - 1 = 3 Change in x values = 2 - 1 = 1 Slope = 3/1 = 3 Set 2 Change in y values = 6 - 7 = -1 Change in x values = 4 - 2 = 2 Slope = -1/2 Compare 3 and -1/2 and you have your answer.

OpenStudy (anonymous):

Thank you